# BuildArena: A Physics‑Aligned Interactive Benchmark of LLMs for Engineering Construction

### v3

@@ Abstract
Engineering construction automation aims to transform natural language specifications into physically viable structures, requiring complex integrated reasoning under strict physical constraints. While modern LLMs possess broad knowledge and strong reasoning capabilities that make them promising candidates for this domain, their construction competencies remain largely unevaluated. To address this gap, we introduce BuildArena, the first physics-aligned interactive benchmark designed for language-driven engineering construction. It contributes to the community in four aspects: (1) a highly customizable benchmarking framework for in-depth comparison and analysis of LLMs; (2) an extendable task design strategy spanning static and dynamic mechanics across multiple difficulty tiers; (3) a 3D Spatial Geometric Computation Library for supporting construction based on language instructions; (4) a baseline LLM agentic workflow that effectively evaluates diverse model capabilities. On eight frontier LLMs, BuildArena comprehensively evaluates their capabilities for language-driven and physics-grounded construction automation. The project page is at build-arena.github.io.
@@ 1 Introduction
Language-driven automated construction presents challenges in two key aspects. On one hand, it necessitates physics simulation environments with high fidelity to real-world constraints so that virtual designs and assembly procedures adhere to geometric, physical, and structural constraints. While modern physics engines and robotics benchmarks offer robust simulation capabilities (Todorov et al., 2012; Coumans & Bai, 2016; Makoviychuk et al., 2021), there remains a gap in environments that integrate physics verification with language-driven multi-component assembly processes. On the other hand, the domain demands multilevel reasoning across long temporal horizons and 3D spatial contexts, as engineering artifacts inherently exhibit hierarchical organization, and their assembly follows sequential dependencies with strict feasibility constraints (Jiménez, 2013; De Fazio & Whitney, 2003). These factors collectively require both breadth of domain knowledge and depth of analytical thinking, challenging even for expert human engineers.
@@
Despite these advances, current evaluations on LLMs provide insufficient evidence of their capacity to construct physical entities. Established LLM benchmarks predominantly assess mathematical and programming capabilities (Cobbe et al., 2021; Hendrycks et al., 2021; Chen et al., 2021), which are evaluated mainly in textual or static environments, without interactions with physical environments. Existing physical reasoning datasets focus on physics understanding, but neglect the multi-step construction processes (Bakhtin et al., 2019; Cherian et al., 2024). Meanwhile, research in programmatic 3D or CAD generation has advanced generation performance but rarely validates whether the generated designs yield executable assemblies under realistic physical conditions (Jones et al., 2020; Mallis et al., 2024). This interdisciplinary gap highlights the absence of frameworks to evaluate whether LLMs can effectively translate natural language specifications into physically viable assemblies. This limitation motivates our research question: How can we comprehensively evaluate LLMs for language-driven and physics‑grounded construction automation?
@@
Benchmarks
Spatial
Reasoning
3D
Construction
Construction-aimed
Planning
Physical
Simulator
Interactive
Environment
@@ 2.1 Task
  • Quantification: Extent of explicit numerical reasoning required (Kamble et al., 2024).
  • Robustness: Tolerance to single-point failures (Zhao et al., 2023).
@@
  • Precision: Strictness of geometric requirement for placement and orientation (Gao et al., 2024b).
  • Ambiguity: Clarity and completeness of task guidance (Moon et al., 2024).
@@ 2.2 Language-driven and Physics-grounded Construction
From the perspective of human engineering practice, construction is inherently an incremental and constraint-driven process. Structures are assembled step by step, each new component must connect to existing ones, and physical feasibility (e.g., collision avoidance) is continuously verified (Wilson & Latombe, 1994; De Fazio & Whitney, 2003). Each successful action requires accurate reasoning about the spatial relationships between new and existing structures. These features necessitate Besiege (Spiderling, 2018), an ideal platform to evaluate the LLMs for physics-grounded construction automation. Besiege is a popular construction sandbox game with realistic physics simulation, widely validated by world-wide player community to align with human physical intuition. It has a rich modules space, a complete collection of basic structural and functional module types that can be combined to build complex objects, all completed by iteratively attachment and refinement of the native modules. Details about modules are provided in Appendix E.1.
@@ 2.3 LLM agentic workflow
LLMs execute the construction procedure via LLM agentic workflow (Zhang et al., 2025a). For clear comparison between different LLMs, we restrict our consideration to workflows where all its entities employ the same LLM, differentiated by their respective prompts, as illustrated in Figure 2. After careful refinements, we finally obtained an effective workflow. Its design follows two principles: (1) a coarse-to-fine structure with an outline plan and progressing to granular, executable details (Xue et al., 2024); (2) a multi-party debate and multi-turn revision framework for incremental improvement of construction quality (Du et al., 2023). Based on these principles, our workflow employs five entities: Planner (P), Drafter (D), Reviewer (R), Builder (B), and Guidance (G). In addition, Controller (C) is used for the task Transport. Prompt details are provided in Appendix G. The workflow includes three stages, as shown in the bottom row of Figure 4:
@@ 2.4 Evaluation
Evaluation metrics cover performance and cost. Performance includes three metrics: (1) Number of parts, referring to the count of modules present in the construction result. A smaller value is preferable because a core principle in engineering prioritizes simpler system structures to improve maintainability and reliability (El-Rayes & Khalafallah, 2005). (2) Success rate, defined as the proportion of trials that successfully passed the criteria among 64 samples, a higher value is better. (3) Performance indicator, a task-specific metric extracted from simulation data that evaluates the performance under realistic physical conditions. A higher value is preferable for all indicators. Detailed success criterion and indicator setup of each task are specified in Section 2.1. Cost is evaluated using three metrics: (1) number of accumulated input tokens, (2) number of output tokens, and (3) total number of LLM requests. A lower value is preferable for all the cost metrics.
@@ Appendix A Related Work
Physics Simulation Environments. Advanced engines (MuJoCo (Todorov et al., 2012), Isaac Gym (Makoviychuk et al., 2021)) and platforms (Autodesk (Autodesk Inc., 2024), SimScale (SimScale GmbH, 2024)) offer robust physics modeling, while digital twins create virtual structure representations (Dai et al., 2024). However, they lack integration with language interfaces for interactive construction, limiting their potential for following human instructions. LLM-based simulation studies focus on decision-making, not mechanical assembly (Kleiman et al., 2025; Gao et al., 2024a), and robot-centric environments (Gazebo, DART-LLM (Wang et al., 2024)) mainly target manipulation rather than structural building. BuildArena uniquely integrates a physics-aligned sandbox with a standardized language interaction protocol, tailored to evaluate the engineering construction driven by LLMs.
AI-Driven Construction Automation. AI applications in construction focus on parameterized design (Newton, 2019) and construction planning optimization (Zhang & Yang, 2025). Integrations with LLMs are still in the early stages(Ma12 et al., ). The critical challenge of translating natural language to physically feasible structural assembly remains to be solved.

### v4

@@ Abstract
Engineering construction automation aims to transform natural language specifications into physically viable structures, requiring complex integrated reasoning under strict physical constraints. While modern LLMs possess broad knowledge and strong reasoning capabilities that make them promising candidates for this domain, their construction competencies remain largely unevaluated. To address this gap, we introduce BuildArena, the first physics-aligned interactive benchmark designed for language-driven engineering construction. It contributes to the community in four aspects: (1) a highly customizable benchmarking framework for in-depth comparison and analysis of LLMs; (2) an extendable task design strategy spanning static and dynamic mechanics across multiple difficulty tiers; (3) a 3D Spatial Geometric Computation Library for supporting construction based on language instructions; (4) a baseline LLM agentic workflow that effectively evaluates diverse model capabilities. On eight frontier LLMs, BuildArena comprehensively evaluates their capabilities for language-driven and physics-grounded construction automation.
@@ 1 Introduction
Language-driven automated construction presents challenges in two key aspects. On one hand, it necessitates physics simulation environments with high fidelity to real-world constraints so that virtual designs and assembly procedures adhere to geometric, physical, and structural constraints. While modern physics engines and robotics benchmarks offer robust simulation capabilities (Todorov et al., 2012; Coumans and Bai, 2016; Makoviychuk et al., 2021), there remains a gap in environments that integrate physics verification with language-driven multi-component assembly processes. On the other hand, the domain demands multilevel reasoning across long temporal horizons and 3D spatial contexts, as engineering artifacts inherently exhibit hierarchical organization, and their assembly follows sequential dependencies with strict feasibility constraints (Jiménez, 2013; De Fazio and Whitney, 2003). These factors collectively require both breadth of domain knowledge and depth of analytical thinking, challenging even for expert human engineers.
@@
Despite these advances, current evaluations on LLMs provide insufficient evidence of their capacity to construct physical entities. Established LLM benchmarks predominantly assess mathematical and programming capabilities (Cobbe et al., 2021; Hendrycks et al., 2021; Chen et al., 2021), which are evaluated mainly in textual or static environments, without interactions with physical environments. Existing physical reasoning datasets focus on physics understanding, but neglect the multi-step construction processes (Bakhtin et al., 2019; Cherian et al., 2024). Meanwhile, research in programmatic 3D or CAD generation has advanced generation performance but rarely validates whether the generated designs yield executable assemblies under realistic physical conditions (Jones et al., 2020; Mallis et al., 2025). This interdisciplinary gap highlights the absence of frameworks to evaluate whether LLMs can effectively translate natural language specifications into physically viable assemblies. This limitation motivates our research question: How can we comprehensively evaluate LLMs for language-driven and physics‑grounded construction automation?
@@
Benchmarks Spatial Reasoning 3D Construction Construction-aimed Planning Physical Simulator Interactive Environment
@@ 2.1 Task
  • Quantification: Extent of explicit numerical reasoning required (Olu-lawal et al., 2024).
  • Robustness: Tolerance to single-point failures (Geng et al., 2025).
@@
  • Precision: Strictness of geometric requirement for placement and orientation (Gang et al., 2024).
  • Ambiguity: Clarity and completeness of task guidance (Moon et al., 2025).
@@ 2.2 Language-driven and Physics-grounded Construction
From the perspective of human engineering practice, construction is inherently an incremental and constraint-driven process. Structures are assembled step by step, each new component must connect to existing ones, and physical feasibility (e.g., collision avoidance) is continuously verified (Wilson and Latombe, 1994; De Fazio and Whitney, 2003). Each successful action requires accurate reasoning about the spatial relationships between new and existing structures. These features necessitate Besiege (Spiderling, 2018), an ideal platform to evaluate the LLMs for physics-grounded construction automation. Besiege is a popular construction sandbox game with realistic physics simulation, widely validated by world-wide player community to align with human physical intuition. It has a rich modules space, a complete collection of basic structural and functional module types that can be combined to build complex objects, all completed by iteratively attachment and refinement of the native modules. Details about modules are provided in Appendix E.1.
@@ 2.3 LLM agentic workflow
LLMs execute the construction procedure via LLM agentic workflow (Zhang et al., 2025a). For clear comparison between different LLMs, we restrict our consideration to workflows where all its entities employ the same LLM, differentiated by their respective prompts, as illustrated in Figure 2. After careful refinements, we finally obtained an effective workflow. Its design follows two principles: (1) a coarse-to-fine structure with an outline plan and progressing to granular, executable details (Xue et al., 2024); (2) a multi-party debate and multi-turn revision framework for incremental improvement of construction quality (Du et al., 2024). Based on these principles, our workflow employs five entities: Planner (P), Drafter (D), Reviewer (R), Builder (B), and Guidance (G). In addition, Controller (C) is used for the task Transport. Prompt details are provided in Appendix G. The workflow includes three stages, as shown in the bottom row of Figure 4:
@@ 2.4 Evaluation
Evaluation metrics cover performance and cost. Performance includes three metrics: (1) Number of parts, referring to the count of modules present in the construction result. A smaller value is preferable because a core principle in engineering prioritizes simpler system structures to improve maintainability and reliability (El-Rayes and Khalafallah, 2005). (2) Success rate, defined as the proportion of trials that successfully passed the criteria among 64 samples, a higher value is better. (3) Performance indicator, a task-specific metric extracted from simulation data that evaluates the performance under realistic physical conditions. A higher value is preferable for all indicators. Detailed success criterion and indicator setup of each task are specified in Section 2.1. Cost is evaluated using three metrics: (1) number of accumulated input tokens, (2) number of output tokens, and (3) total number of LLM requests. A lower value is preferable for all the cost metrics.
@@ Appendix A Related Work
Physics Simulation Environments. Advanced engines (MuJoCo (Todorov et al., 2012), Isaac Gym (Makoviychuk et al., 2021)) and platforms (Autodesk (Autodesk Inc., 2024), SimScale (SimScale GmbH, 2024)) offer robust physics modeling, while digital twins create virtual structure representations (Dai et al., 2024). However, they lack integration with language interfaces for interactive construction, limiting their potential for following human instructions. LLM-based simulation studies focus on decision-making, not mechanical assembly (Kleiman et al., 2025; Gao et al., 2024), and robot-centric environments (Gazebo, DART-LLM (Wang et al., 2024)) mainly target manipulation rather than structural building. BuildArena uniquely integrates a physics-aligned sandbox with a standardized language interaction protocol, tailored to evaluate the engineering construction driven by LLMs.
AI-Driven Construction Automation. AI applications in construction focus on parameterized design (Newton, 2019) and construction planning optimization (Zhang and Yang, 2025). Integrations with LLMs are still in the early stages(Ma12 et al., ). The critical challenge of translating natural language to physically feasible structural assembly remains to be solved.